What temperature would a 2m water solution of CaCl2 boil at?

Kb water is .52 BP water is 100oC ATb = (i)(kb)(m)
1
Expert's answer
2020-04-03T11:49:42-0400

ΔT=Kb*m= 0.52*2=1.04 oC;

Answer: a 2m water solution of CaCl2 would boil at 101.04 oC